# 4.1. Degenerate and non-degenerate components Let us use the notation in (3.3) for the Littlewood-Paley partition of unity. Typically, runs over the set of integers, but also we are going to consider the functions where runs over the set of integers and is fixed, and then In terms of the projective co-ordinates in (3.1), in view of the homogeneity of the phase function as defined in (4.2), we can write when In such a case we are going to simplify the notation by writing

### v4

@@ 4.4. Control of the error operator
In this subsection we prove that the error operator is bounded on The idea behind the proof of this fact can be understood from the calculus of Fourier integral operators. The operator is essentially a Fourier integral operator with the same phase function that in the sense that its phase evaluated at the stationary point is see [46, Page 19, Eq. (16)]. So, intuitively, if would have had the form in (1.1), namely, one could compute an asymptotic expansion for the symbol of as follows
@@
for any
Below we will prove that essentially and have the same principal symbol and then the order of the error operator should be, at least heuristically, the order of the ’s when that is This order is good enough to deduce the -boundedness of However, since the definition of also involves integration with respect to the structure of this operator is not as the one in (1.1), and one has to use the stationary phase method to prove that each dyadic component of the error has a kernel with a -norm satisfying uniformly in a decay proportional to Then, if proved this we are allowed to use Proposition 6.1 in [46, Page 17] to deduce the boundedness of even under the presence of the oscillating term
@@
By following the argument in [46, Page 19] one can make the approximation
@@
Here, the main integral to study is Indeed, arguing as in [46, Page 17] one can demonstrate that is a smoothing symbol. Indeed, one can estimate this integral over the region by inserting the cut-off in the integrand of Since on and in view of the identity
@@
Note that the integration with respect to is over the region and then we can estimate this integral by introducing a cutt-off
In accordance with the principle of stationary phase, we now look at where the phase is stationary in From the homogeneity of the phase function we have that
@@
Now, we have prepared the way in order to use Proposition 6.1 of [46] for deducing the estimate (4.55). Indeed, our estimate in (4.55) is the analog of (19) in [46, Page 19] which is a consequence of [46, Proposition 6.1]. For completeness, we present such a proposition as follows.
@@
So, note that for the validity of (4.55)
we have used Lemma 4.13 applied to The proof of the -boundedness of is complete.
@@
With the notation in Subsection 4.1 the operator has been decomposed into its degenerate and non-degenerate components, and respectively. In Subsection 4.2 in Proposition 4.4 we have proved that the degenerate operator is bounded on As for the non-degenerate operator we have constructed in Subsubsection 4.3.1 a Fourier integral operator which is an average operator, of the same order that and a pseudo-differential operator of order zero. We have proved, that the factorisation approach introduced in Tao [46] also operates in our case, namely, that is bounded on (see Proposition 4.9) is of weak (1,1) type in view of the standard Calderón-Zygmund theory, and that the error operator is bounded on see Subsection 4.4. Since then is of weak (1,1) type. The proof of Theorem 3.1 is complete.

### v5

# The weak (1,1) boundedness of Fourier integral Operators with complex phases

@@ 4.4. Control of the error operator
In this subsection we prove that the error operator is bounded on The idea behind the proof of this fact can be understood from the calculus of Fourier integral operators. The operator is essentially a Fourier integral operator with the same phase function that in the sense that its phase evaluated at the stationary point is see [47, Page 19, Eq. (16)]. So, intuitively, if would have had the form in (1.1), namely, one could compute an asymptotic expansion for the symbol of as follows
@@
for any Below we will prove that essentially and have the same principal symbol and then the order of the error operator should be, at least heuristically, the order of the ’s when that is This order is good enough to deduce the -boundedness of However, since the definition of also involves integration with respect to the structure of this operator is not as the one in (1.1), and one has to use the stationary phase method to prove that each dyadic component of the error has a kernel with a -norm satisfying uniformly in a decay proportional to Then, if proved this we are allowed to use Proposition 6.1 in [47, Page 17] to deduce the boundedness of even under the presence of the oscillating term
@@
By following the argument in [47, Page 19] one can make the approximation
@@
Here, the main integral to study is Indeed, arguing as in [47, Page 17] one can demonstrate that is a smoothing symbol. Indeed, one can estimate this integral over the region by inserting the cut-off in the integrand of Since on and in view of the identity
@@
Note that the integration with respect to is over the region and then we can estimate this integral by introducing a cutt-off In accordance with the principle of stationary phase, we now look at where the phase is stationary in From the homogeneity of the phase function we have that
@@
Now, we have prepared the way in order to use Proposition 6.1 of [47] for deducing the estimate (4.55). Indeed, our estimate in (4.55) is the analog of (19) in [47, Page 19] which is a consequence of [47, Proposition 6.1]. For completeness, we present such a proposition as follows.
@@
So, note that for the validity of (4.55) we have used Lemma 4.13 applied to The proof of the -boundedness of is complete.
@@
With the notation in Subsection 4.1 the operator has been decomposed into its degenerate and non-degenerate components, and respectively. In Subsection 4.2 in Proposition 4.4 we have proved that the degenerate operator is bounded on As for the non-degenerate operator we have constructed in Subsubsection 4.3.1 a Fourier integral operator which is an average operator, of the same order that and a pseudo-differential operator of order zero. We have proved, that the factorisation approach introduced in Tao [47] also operates in our case, namely, that is bounded on (see Proposition 4.9) is of weak (1,1) type in view of the standard Calderón-Zygmund theory, and that the error operator is bounded on see Subsection 4.4. Since then is of weak (1,1) type. The proof of Theorem 3.1 is complete. ∎
5. Final remarks
5.1. Applications
In order to illustrate the applications of Theorem 3.1 to a priori estimates for the Cauchy problem with complex characteristics we consider the case where the associated operator is a classical pseudo-differential operator of order one. Let us introduce some required notation. Let be a compact Riemmanian manifold of dimension and let us consider Denote by the positive Laplacian associated to the metric Denote the space defined by the norm
We also consider the weak -space defined by the seminorm
and by its local version, namely the family of functions such that for any compactly supported function
Denote the partial derivative with respect to the time variable and let Let us consider the Cauchy problem for a first-order classical pseudo-differential operator
We assume that its principal symbol has non-negative imaginary part, namely, satisfies
The solution operator of (5.1) can be realised as a Fourier integral operator of order zero with a global complex phase function of positive type, see [40, Page 117] and also Laptev, Safarov and Vassiliev [30] for details. Applying (essentially) Theorem 3.1 to the operator
of order we have that for any compactly supported initial datum , and for a fixed time the solution of the Cauchy problem (5.1) satisfies This analysis proves the following consequence of Theorem 3.1.
Corollary 5.1.
Let and let be compactly supported. Then, for a fixed , the solution of the Cauchy problem (5.1) satisfies
Proof.
We write namely, the solution of (5.1) in terms of the solution operator Since we have that
We can use the factorisation
and the fact that is a Fourier integral operator with complex phase of order 333See [40, Page 117] and also Laptev, Safarov and Vassiliev [30] for details. to prove that Indeed, if denotes a partition of unity subordinated to a finite open covering of the compact manifold then and We can assume that the covering is an atlas for the manifold and then each is diffeomorphic to an Euclidean open set Note that is a Fourier integral operator of order that can be microlocalised to a Fourier integral operator whose symbol is compactly supported in the spatial variable. Here, denotes the multiplication operator by the function Then, Theorem 3.1 imply that for any we have that
which proves Corollary 5.1. ∎
We also observe that a similar analysis can be adapted to investigate the mapping properties of the parametrix of other first-order complex partial differential operators, see e.g. [40, Page 120] for a problem that appears naturally in the study of the oblique derivative problem. This was observed e.g. by Melin and Sjöstrand in [31]. For other applications of the theory of Fourier integral operators with complex phases to a priori estimates for the Cauchy problem with complex characteristics we refer to Trèves [48, Chapter XI] and to [40, Chapter 5].
5.2. Conclusions and bibliographical discussion
In this work we have analysed the weak (1,1) boundedness of Fourier integral operators with complex-valued phase functions. As in the setting of real-valued phase functions regarding the weak (1,1) boundedness due to Tao in [47], our estimate is local in the sense that we have considered the symbol to be compactly supported in the -variable, following the previous approaches, see Seeger, Sogge and Stein [45], Beals [3], and the monograph of the second author about the mapping properties of Fourier integral operators with complex phases [40]. Among other things, the sharpness of the order for the weak (1,1) inequality of elliptic Fourier integral operators has been discussed by the authors in [9], and we also refer to [39] for the sharpness of Seeger-Sogge-Stein orders regarding the -boundeness of these operators.
There has been in the last decades a wide interest in the analysis of the global continuity properties for Fourier integral operators. This problem can be traced back to the work of Asada and Fujiwara [2] and of Peral [36] and Miyachi [34]. We refer to [41], [42] and [12, 13] for global properties on -spaces. As for global criteria for Fourier integral operators on Hörmander classes removing the calculus condition and allowing the ranges and we refer the reader to Dos Santos Ferreira and Staubach [14], Kenig and Staubach [28] and the works of Staubach and his collaborators [26, 27, 37, 38]. We also refer to the seminal work of Álvarez and Hounie [1] for the boundedness of pseudo-differential operators without the condition which was the starting point for the aforementioned results in the setting of Fourier integral operators. For the global parametrization of Lagrangian distributions we refer the reader to Laptev, Safarov, and Vassiliev [30] and to [40]. We also refer to the work [10] for the global definition of Fourier integral operators on compact Lie groups in terms of the group Fourier transform.
The theory of Fourier integral operators with complex phases was developed by Melin and Sjöstrand in [31] motivated by the problem of the construction of parametrices for operators of principal type with non-real principal symbols, see for instance [32]. The singularities of the Bergman kernel can be also approximated in terms of asymptotic expansions of certain Fourier integral operators with complex-valued functions, see Boutet de Monvel and Sjöstrand [7]. For a systematic analysis of the Bergman kernel we refer to Fefferman [18, 19].
There has been also wide activity concerning the smoothing estimates for Fourier integral operators after the proof of Bourgain and Demeter [6] of the decoupling conjecture. The smoothing effect for Fourier integral operators was conjectured by Sogge in [46]. It will be difficult to review the literature about this problem here, but in what follows we share some works in this direction. Indeed, we refer the reader to Bourgain [5], Minicozzi and Sogge [33], Guth, Wang and Zhang [21], Beltran, Hickman and Sogge [4], to Gao, Liu, Miao, and Xi [20] and references therein for details.
Conflict of interests statement - Data Availability Statements The authors state that there is no conflict of interest. Data sharing does not apply to this article as no datasets were generated or analysed during the current study.
5.3. Acknowledgement
The authors thank Terence Tao for suggesting the factorisation approach used in this manuscript.